OpenStudy (anonymous):

A right circular cylinder has a base whose area is 616 square feet. If the height of the cylinder is 13 feet, what is its lateral surface area? Use π = 22 over 7 Answer 363 ft2 572 ft2 1144 ft2 1760 ft2

OpenStudy (anonymous):

Get the radius from: \[Area_{Circle} = \pi r^2\] and fidn the lateral surface area from un wrapping the cylinder like such: 

OpenStudy (anonymous):

but doesnt r= radius? how do i work that out if i don't have r forgive meif i sound blonde right now

OpenStudy (anonymous):

math isnt my subject.. at all

OpenStudy (anonymous):

r is indeed the radius. you aren't given the radius but you are given the area which you can use to find the radius. try reorganizing the equation i posted so that you end up with 'r' by itself

OpenStudy (anonymous):

i figured i had to use the are to find the radius but idk how to do that with the equation.. i dont think ill look at it agaiin

OpenStudy (anonymous):

unless i can just half it but that would mean the givin number is the diameter and it says area not diameter..

OpenStudy (anonymous):

Take what you know and sub it in. Then you'll have to rearrange until you've got what your looking for by it's lonesome. In this case you have an equation of two variables, if you are given one you can solve the other. Given the area we can proceed: \[Area_{circle} = 616 = \pi r^2\] \[616 = \pi r^2\] \[616/\pi = r^2\] \[\sqrt{616/\pi} = r\]

OpenStudy (anonymous):

ok i got 14.0 .. i devided 616 by 3.14 and then found the square root of that nd that what i got

OpenStudy (anonymous):

Good, so that's your radius. Now the 'lateral area' of the cylinder represents the side area. if you can imagine unpeeling a soup can label then it's not a stretch to say that this round area becomes just a rectangle(see the picture above).

OpenStudy (anonymous):

yeahh so now just plug in my radius into the equation and thats my answer?